Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Combinatorics: Solution Explanation

Expert replies
by Dblooos » Sat Dec 28, 2013 8:35 am
A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The solution says:

None of the 14 patients is "special" in any way, so each of them must have the same
chance of receiving Progaine or Ropecia. Since Progaine is only administered to one patient,
each patient (including Donald) must have probability 1/14 of receiving it. The same logic
also holds for Ropecia. Since Donald cannot receive both of the medicines, the desired
probability is the probability of receiving Progaine, plus the probability of receiving Ropecia:
1/14 + 1/14 = 1/7.

Is this correct? If yes how? If not what should be the correct answer? Please help. Thank you.
Join the discussion
Source: — Problem Solving |

by theCodeToGMAT » Sat Dec 28, 2013 8:39 am
14 --> 1(Progaine)
13 --> 1(Ropecia)
12 --> 1

To find: Prob of Progaine or Ropecia

Probability of (Progaine) = 1/14

Probability of (Ropecia) = 13/14*1/13 = 1/14

Probability = 1/14 + 1/14 = 1/7
R A H U L
Join the discussion

by Stuart@KaplanGMAT » Sat Dec 28, 2013 8:46 am
Dblooos wrote:A medical researcher must choose one of 14 patients to receive an experimental
medicine called Progaine. The researcher must then choose one of
the remaining 13 patients to receive another medicine, called Ropecia.
Finally, the researcher administers a placebo to one of the remaining 12
patients. All choices are equally random. If Donald is one of the 14 patients,
what is the probability that Donald receives either Progaine or Ropecia?

The solution says:

None of the 14 patients is "special" in any way, so each of them must have the same
chance of receiving Progaine or Ropecia. Since Progaine is only administered to one patient,
each patient (including Donald) must have probability 1/14 of receiving it. The same logic
also holds for Ropecia. Since Donald cannot receive both of the medicines, the desired
probability is the probability of receiving Progaine, plus the probability of receiving Ropecia:
1/14 + 1/14 = 1/7.

Is this correct? If yes how? If not what should be the correct answer? Please help. Thank you.
Hi Dbloos!

That answer is in fact correct.

Here's another way of looking at it: 2 out of the 14 patients will be chosen to receive P or R. Since the selection is random, each patient has a 2/14 chance of being one of the two.

We can confirm this answer with common sense (a very powerful and often under-utilized tool on the GMAT!): if Donald didn't have a 2/14 chance of getting exactly 1 of those two treatments, then at least one of the other 14 applicants would have to have a greater than 2/14 chance to get the probabilities to properly add up. Since the choices are completely random, that makes no sense!

A big takeaway from this question - and something that stumps a lot of people on probability questions - is that when you make selections without replacement, it's irrelevant whether you make them simultaneously or sequentially. So, if the question had said that all 3 patients were chosen at the same time, the answer would remain the same (2/14 for getting either P or R; 3/14 for getting either P, R or the placebo).
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by Dblooos » Sat Dec 28, 2013 8:56 am
Thank you Rahul and Stuart. It makes sense now ;)
Join the discussion

by Brent@GMATPrepNow » Sat Dec 28, 2013 9:09 am
Dblooos wrote:A medical researcher must choose one of 14 patients to receive an experimental medicine called Progaine. The researcher must then choose one of the remaining 13 patients to receive another medicine, called Ropecia. Finally, the researcher administers a placebo to one of the remaining 12 patients. All choices are equally random. If Donald is one of the 14 patients, what is the probability that Donald receives either Progaine or Ropecia?
We can also solve this using the complement.

P(Donald receives either Progaine or Ropecia) = 1 - P(Donald receives neither Progaine nor Ropecia)

P(Donald receives neither Progaine nor Ropecia)
At this point, I like to ask, " What exactly must occur in order for this event to happen?"
Well, for Donald to receive neither Progaine nor Ropecia it must be the case that he does not receive Prograine during the first selection and he does not receive Ropecia during the second selection.

In other words, . . .
P(Donald receives neither Progaine nor Ropecia) = P(Donald does not receive Prograine during the 1st selection AND he does not receive Ropecia during the 2nd selection.
= P(Donald does not receive Prograine during the 1st selection) x P(Donald does not receive Ropecia during the 2nd selection)
= (13/14)x(12/13)
= 12/14
= 6/7

So, P(Donald receives either Progaine or Ropecia) = 1 - (6/7)
= [spoiler]1/7[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion